OpenStudy (anonymous):

find the standard deviation of 18,19,20,21,22

OpenStudy (the_fizicx99):

1. Find the mean: \(\ \dfrac{18 + 19 + 20 + 21 + 22}{5} \) Mean = 20 2. subtract the original value with the mean, 18 - 20 = -2 19 - 20 = -1 20 - 20 = 0 21 - 20 = 1 22 - 20 = 2 Now square that, \(\ -2^2 = 4 \) \(\ -1^2 = 1\) \(\ 0^2 = 0\) \(\ 1^2 = 1\) \(\ 2^2 = 4\) Now divide by add them up and divide by 5, \(\ \dfrac{ 4 + 1 + 0 + 1 +4}{5} = \dfrac{10}{5} = 2\) \(\sigma = \sqrt{2} = 1.41 \)

OpenStudy (the_fizicx99):

The original formula is: \[\sigma =\sqrt{\Sigma\dfrac{(x - mean)^2}{n}}\]

OpenStudy (kc_kennylau):

Mean = Sum / number Variance = Mean of distances to mean squared S.D. = square root of variance